OpenStudy (anonymous):

A concrete mix is to have mass proportions of cement: fine aggregates: coarse aggregates of 1.0: 1.9: 2.8, with a water-cement ratio of 26.5 litres of water per 42.6 kg of cement. The unit mass of each of these mix components is given below.

OpenStudy (anonymous):



OpenStudy (anonymous):

Determine the mass of each component required to produce 38.0 m3 (38,000 litres) of concrete. Use the notation: C = mass of cement (kg), F = mass of fine aggregates (kg), G = mass of coarse aggregates (kg), and W = mass of water (kg). Solve for C, F, G, and W using Gauss elimination.

OpenStudy (anonymous):

@ganeshie8 Could you help for for a bit? I just don't get how to set up the matrix.

OpenStudy (anonymous):

@ParthKohli or @thomaster Could you help if you have time?

OpenStudy (anonymous):

@Compassionate or @dan815 Do you know how to set this up?

OpenStudy (anonymous):

Hint: a ratio between two components provides one linear equation, e.g. C/F = 1.0/1.9, which gives the linear equation: 1.90C – 1.0F = 0 (the coefficients for G and W are zero in this equation).